OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1 3x + 2y = -4 x + 2y = -8 5x - 4y = 14 x - 2y = -2 x + 3y = -4 4x + 5y = 5 x + 3y = -5 2x - y = 18

OpenStudy (ivettef365):

y=-4x-1 now substitute (-4x-1) for y on the equation: 6x+2y=0

OpenStudy (ivettef365):

You have 6x + (2(-4x-1)) = 0

OpenStudy (ivettef365):

Then 6x - 8x - 2 = 0

OpenStudy (ivettef365):

Therefore -2x = 2 ---- --- -2 -2 x = -1

OpenStudy (anonymous):

i dont understand substitution

OpenStudy (ivettef365):

Then to find y substitute -1 for x 6x + 2y = 0 6(-1) + 2y = 0 -6 + 2y = 0 2y = 6 y = 3

OpenStudy (ivettef365):

Therefore x = -1 and y = 3 Now you can check by substitution of x and y to see if is true

OpenStudy (ivettef365):

6x + 2y = 0 6(-1) + 2(3) = 0 -6 +6 = 0 0=0 this is true

OpenStudy (ivettef365):

The other equation: 4x + y = -1 4(-1) + 3 = -1 -4 + 3 = -1 -1 = -1 true then is correct

OpenStudy (anonymous):

@ivettef365 what is x + 2y = -8 ?

OpenStudy (anonymous):

?

OpenStudy (ivettef365):

on this one I would live x by itself x = -2x-8

OpenStudy (ivettef365):

Now substitute x by (-2x-8) on equation 3x + 2y = -4

OpenStudy (anonymous):

3(-2x-8)+2y=-4?

OpenStudy (ivettef365):

Sorry I meant x = -2y - 8

OpenStudy (ivettef365):

2(-2y-8) + 2y = -4

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

then just solve it? @ivettef365

OpenStudy (ivettef365):

right, if you get stuck let me know

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

@ivettef365 2(-2y-8) + 2y = -4 -4y - (-16) + 2y = -4 -4y+16 + 2y = -4 -2y +16 = -4 -2y = -20 y=-10 ?

OpenStudy (anonymous):

then: 3x + 2y = -4 3x + 2(-10) = -4 3x+-20=-4 3x=-24 x=-8 ? Is this right for number two?

OpenStudy (ivettef365):

I got for x = 2 and for y = -5

OpenStudy (ivettef365):

You put 2 instead of 3, it should have been 3(-2y-8)+2y = -4

OpenStudy (ivettef365):

-6y-24+2y=-4 -4y-24 = -4 -4y = 20 y = -5

OpenStudy (ivettef365):

Now that you know y = -5 find x use the most simple one x - 2y = -8 x -2(-5) = -8 x + 10 = -8 x = 2

OpenStudy (anonymous):

oohh but the first part is right?

OpenStudy (ivettef365):

Try doing the 3rd problem

OpenStudy (ivettef365):

x = 2y - 2

OpenStudy (ivettef365):

Substitute (2y-2) for x on 5x - 4y = 14

OpenStudy (anonymous):

Alright! x - 2y = -2 x=2y-2 5(2y-2) - 4y = 14

OpenStudy (anonymous):

10y-10 - 4y=14?

OpenStudy (anonymous):

10y+10-4y=14 *

OpenStudy (anonymous):

14y +10=14 14y=4 y=3.5? @ivettef365

OpenStudy (ivettef365):

10y-4y-10 = 14 6y = 14 +10 6y = 24 y= 4

OpenStudy (anonymous):

mind if i work the other two with you?

OpenStudy (anonymous):

5x - 4(4) = 14 5x-16=14 5x=30 x=6

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

Now to check you can do 5(6) - 4(4) = 14 30 - 16 = 14 14 = 14 correct

OpenStudy (ivettef365):

The other one x - 2y = -2 6 - 2(4) = -2 6 - 8 = -2 -2 = -2 correct

OpenStudy (anonymous):

4x + 5y = 5 5y+4x=5?

OpenStudy (anonymous):

I need help with the second equation for both problems :/ I'm not sure how to convert them

OpenStudy (ivettef365):

This one you use x + 3y = -4

OpenStudy (ivettef365):

Convert to x = -3y-4

OpenStudy (anonymous):

wheres the 5...?

OpenStudy (anonymous):

err for 4x + 5y = 5. i thought we convert the second equation first

OpenStudy (ivettef365):

It doesn't matter which one you convert, use the easiest one to convert

OpenStudy (ivettef365):

Now substitute (-3y-4) for x on equation 4x + 5y = 5

OpenStudy (anonymous):

do you mean for y...?

OpenStudy (ivettef365):

For x

OpenStudy (ivettef365):

You do the following 4(-3y-4) + 5y = 5

OpenStudy (anonymous):

4(-3y-4)+5y=5 -12y +5y = 5

OpenStudy (anonymous):

-7y = 5

OpenStudy (ivettef365):

-12y -16+5y = 5

OpenStudy (ivettef365):

You missed the -16

OpenStudy (anonymous):

where is the -16 come from? i thought it was 4*-4

OpenStudy (ivettef365):

4(-3y-4)+5y = 5 -12y-16+5y = 5 (you multiply the 4 by -3y and by -4)

OpenStudy (ivettef365):

4*-4 = -16

OpenStudy (anonymous):

-12y -16+5y = 5 -12y+5y-16=5 -7y-16=5 -7y=21 y=-3?

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

yay! one more ^_^

OpenStudy (ivettef365):

Now solve for x

OpenStudy (anonymous):

i keep forgetting that! 4x + 5y = 5 4x+5(-3) = 5 4x+-15=5

OpenStudy (anonymous):

or -15+4x=5?

OpenStudy (ivettef365):

When you already have a value, like in this case you have the value for y = -3 try to use the easiest equation to find the other value

OpenStudy (ivettef365):

In this case is x + 3y = -4 x + 3(-3) = -4

OpenStudy (anonymous):

x + 3y = -4 x+3(-3-- you beat me to it :P

OpenStudy (anonymous):

x + 3(-3) = -4 x+-15=-4

OpenStudy (ivettef365):

x - 9 = -4

OpenStudy (ivettef365):

+3 * -3 = -9

OpenStudy (anonymous):

yeah.. my calculator said -15? :|

OpenStudy (anonymous):

x-9=-4 x=5

OpenStudy (ivettef365):

correct, then you have x = 5 and y = -3 you can do the checking if you want to

OpenStudy (anonymous):

one more question :D i need you to convert it for me xD

OpenStudy (ivettef365):

The top one is easier to convert

OpenStudy (ivettef365):

You have x + 3y = -5

OpenStudy (ivettef365):

need to transfer the 3y to the other side so the x stays by itself, therefore to transfer you need to switch sign.

OpenStudy (anonymous):

3y-x=-5..?

OpenStudy (ivettef365):

It would be as follows: x + 3y = -5 x = -3y -5

OpenStudy (anonymous):

yeah s: i'm not good at that... 2(-3y -5) - y = 18

OpenStudy (ivettef365):

That is good, what is the next step

OpenStudy (anonymous):

-6y - (-10) - y = 18

OpenStudy (ivettef365):

-6y-10-y = 18

OpenStudy (anonymous):

-6y-y-10=18

OpenStudy (anonymous):

-7y-10=18

OpenStudy (anonymous):

-7y=28 y=-4

OpenStudy (ivettef365):

Good job

OpenStudy (anonymous):

x + 3y = -5 x + 3(-4) = -5

OpenStudy (anonymous):

x+-12=-5

OpenStudy (ivettef365):

did you get the answer

OpenStudy (anonymous):

i was making sure that was right, haha.

OpenStudy (anonymous):

x+-12=-5 x=7?

OpenStudy (ivettef365):

Good job

OpenStudy (anonymous):

thank you so much!!

OpenStudy (ivettef365):

yw :)